OpenStudy (anonymous):

what is ??? ((1/2)-1)/2

OpenStudy (anonymous):

\[((1\div2)-1)\div2\]

OpenStudy (binary3i):

-1/4

OpenStudy (anonymous):

what i think it's -1?

OpenStudy (hoblos):

((1/2)-1)/2 = (-1/2)/2 = -1/4

jhonyy9 (jhonyy9):

-1/2/2 = -1/4